A window glass has dimensions, 1.18 m high, 2.19 mPlease show working. Thanks
Answer #1

given that

high h=1.18 m

wide b=2.19 m

thickness L=1.448 cm=1.448*10^-2 m

higher temperature T2=32 c

lower temperature T1=24 c

co efficient of thermal conductivity K=0.81wm^-1 C^-1

area of the glass A=1.18*2.19=2.5842 m^2

this problem belongs to principle of conduction

according to heat produced in conduction

thermal coductivity = Q=kA(T2-T1) t/l

where Q= heat flow through the glass

k=co.efficient of thermal conductivity

A=area of glass

(T2-T1) = temperature difference

t= time

L=thickness of glass

so the rate of heat follow =P=Q/t = kA(T2-T1)/L

=0.81*2.5842*(32-24)/1.448*10^-2

=1156.5

=1156.5 watts

therefore the rate of heat transfer P=1156.5 watts